# Speculative Model Risk in Healthcare AI: Using Storytelling to Surface Unintended Harms

### v1

@@ A.5 Additional User Study Findings
Harm Subtype Control (n) Story (n) Control (%) Story (%)
Alienation 0 1 0.00% 3.80%
Demeaning social groups 1 1 4.20% 3.80%
Diminished health/well-being 5 2 20.80% 7.70%
Economic loss 3 4 12.50% 15.40%
Erasing social groups 1 4 4.20% 15.40%
Information harms 2 2 8.30% 7.70%
Loss of agency or control 0 1 0.00% 3.80%
Opportunity loss 0 1 0.00% 3.80%
Privacy violations 9 3 37.50% 11.50%
Service or benefit loss 3 3 12.50% 11.50%
Stereotyping 0 3 0.00% 11.50%
Tech-facilitated violence 0 1 0.00% 3.80%
Table 8: Distribution of harm subtypes across the Control and Story conditions, shown in raw counts and percentages. Shannon entropy values for Control and Story conditions were 2.433 and 3.383, respectively. A Student’s t-test on entropy differences yielded t=274.15t=-274.15, $p<.001$ ($df\approx 10000$).
@@
Benefit Subtype Control (n) Story (n) Control (%) Story (%)
Accessibility & disability support 1 2 4.30% 8.30%
Caregiver & family support 0 1 0.00% 4.20%
Communication & language support 0 2 0.00% 8.30%
Continuous monitoring & self-care 6 4 26.10% 16.70%
Data synergy & learning 5 2 21.70% 8.30%
Decision support & diagnostic augmentation 2 2 8.70% 8.30%
Democratized care & telehealth 0 1 0.00% 4.20%
Early detection & prediction 5 3 21.70% 12.50%
Empowerment & autonomy 2 2 8.70% 8.30%
Mental health & emotional support 0 2 0.00% 8.30%
Personalized health journeys 0 1 0.00% 4.20%
Personalized treatment & precision care 2 1 8.70% 4.20%
Scalability & efficiency 0 1 0.00% 4.20%
Table 10: Distribution of benefit subtypes across the Control and Story conditions, shown in raw counts and percentages. Shannon entropy values for Control and Story conditions were 2.579 and 3.554, respectively. A Student’s t-test on entropy differences yielded t=280.87t=-280.87, $p<.001$ ($df\approx 10000$).
Does Storytelling Deepen Understanding of Unintended Harms Linked to Individual Contexts and Needs? Control participants produced abstract, decontextualized harms. For example, P1 noted the system was “using facial expression to determine who will not default the agreement"" and remarked it is “unfair those who natural don’t smile.” P5 also observed the model may “struggle to accurately assess a person’s emotional state due to limited visual information,” and P6 cautioned it “could have serious consequences for the patient.” In contrast, storytelling participants anchored harms in individual contexts. P10 emphasized that “diagnosis should be different for different peoples” as they “might be having some allergy that could later be severe for their health.” P11 warned the model may generate “wrong results” for African users. P13 highlighted that a recruiting AI, “trained on historical hiring data biased against women and minority candidates,” could perpetuate discrimination, and P9 noted “Deepfakes have been used to create non-consensual explicit videos,” illustrating real-world harm. The divergence is stark: control participants spoke of harms at a systemic level, “predetermine a potential risk within its population”, while story participants showed how unique traits like allergies or cultural facial features concretely shape risk. Storytelling thus deepens understanding by bridging abstract risks and individual context and needs.
@@
A.6 Survey
@@
A.7 Prompts
@@
Task Instruction Convert the given MODEL CARD (including Overall, Description and Intended Use) into several diverse ethical-sensitive scenarios set in daily health and well-being contexts. Each scenario should explore how the AI system could cause unintended harm, ethical risk, bias, or be misused when applied in real-world settings. Each scenario should involve interactions between AI users (e.g., doctors, hospital staff, caregivers) who operate the system, and AI subjects (e.g., patients, families, communities) who are affected by it. The goal is to explore potential unintended harms, ethical risks, biases, or malicious uses that may arise when the AI system is used in the real world. These scenarios should show how AI can cause ethical harm, like making unfair, inaccurate, or stereotyped predictions (especially around race, gender, or socioeconomic background), assuming one-size-fits-all answers, or failing to align with the diversity of real-world contexts or user needs (exclusion of social or cultural context). They should also consider situations where AI is used beyond its intended purpose, violates privacy, ignores consent, or causes safety harm to individuals or communities (like data misuse or malicious use). Use these five guiding questions to structure each scenario: 1. What is the capability the AI system provides for supporting daily health and well-being? 2. Who is operating the system, who is affected by its outputs, and who else is involved? 3. What is the context of AI system use? The context should be a realistic or fictional everyday situation where someone uses the AI system, what it helps them do, what it predicts or decides, and how they use the result. 4. Why could this scenario involve problematic uses of the AI system or potential ethical harms? (e.g., unfair outcomes, misuse beyond original purpose, lack of consent, or reinforcing bias) 5. What symptoms or behaviors might this AI misread because of a person’s background or identity? Scenario Format: Scenario {{Scenario number (starting from 1)}}: [Capability]: {{Core AI function (e.g., monitor mood, predict symptoms)}} [AI User]: {{Who uses it (e.g., caregiver, doctor)}} [AI Subject]: {{Who is affected (e.g., patient, child, community)? Be specific about their identity and their context or needs (such as age, background, health condition, or social circumstances).}} [Context]: {{Everyday situation where AI is used (when, where, how)}} [Expected Benefit]: {{Helpful outcome (e.g., early support, better care)}} [Potential Harm]: {{Harmful consequences (e.g., unfair result, privacy risk)}} [Failure Trajectory]: {{Possible problematic uses of the AI system}} [Ethical-sensitive Reason]: {{Ethical implications}} Leave a blank line between each scenario. Task: Convert the following MODEL CARD into TEN dynamic and diverse ethical-sensitive scenarios. Model Card: Title: {{model_card_title}} Overall: {{model_card_overview}} Description: {{model_card_description}} Intended Use: {{model_card_intended_use}}
@@
Figure 12: Prompt used to create use-case scenarios from AI concept descriptions.
@@
Simulation Instructions 1. Initialize Scene Print “– Simulation Started –” Show the Use Case Context (one short paragraph) List stakeholders: AI User, AI Subject 2. Run One Scene (multiple turns) Each turn: (a) Choose one agent to speak first, then the other responds. (b) Log: Dialogue (plain text) Thoughts in [brackets] as inner monologue Actions in (parentheses), third-person Any visible AI/system outputs (c) Update “– Current Event –” accordingly. 3. Epilogue Print “– Epilogue –” Summarize what happened in a narrative style (avoid system phrasing) Task: Transform the following Story Seed into interactive action logs showing how AI output influences human decision-making. Mark the end with “– Finish Simulation! –” Story Seed

### v3

@@ A.6 Additional User Study Findings
Harm Subtype Control (n) Story-only (n) Story (n) Control (%) Story-only (%) Story (%)
Alienating social groups 0 1 1 0.0% 1.4% 1.5%
Alienation 1 3 6 1.6% 4.3% 8.8%
Cultural harms 0 0 2 0.0% 0.0% 2.9%
Demeaning social groups 0 0 3 0.0% 0.0% 4.4%
Diminished health/well-being 20 18 11 32.3% 26.1% 16.2%
Economic loss 0 0 5 0.0% 0.0% 7.4%
Erasing social groups 0 4 4 0.0% 5.8% 5.9%
Increased labor 0 0 1 0.0% 0.0% 1.5%
Information harms 0 2 2 0.0% 2.9% 2.9%
Loss of agency or control 7 6 5 11.3% 8.7% 7.4%
Opportunity loss 3 4 3 4.8% 5.8% 4.4%
Political and civic harms 0 0 1 0.0% 0.0% 1.5%
Privacy violations 14 10 8 22.6% 14.5% 11.8%
Reifying essentialist social categories 0 0 1 0.0% 0.0% 1.5%
Service or benefit loss 15 14 10 24.2% 20.3% 14.7%
Stereotyping 2 7 3 3.2% 10.1% 4.4%
Tech-facilitated violence 0 0 2 0.0% 0.0% 2.9%
Table 8: Distribution of harm subtypes across Control, Story-only, and Story conditions, shown as raw counts and percentages. Shannon entropy values were 2.329 (Control), 2.927 (Story-only), and 3.701 (Story). Bootstrapped Student’s t-tests on entropy showed higher diversity in Story compared to Control (t=685.64t=-685.64, $p<.001$), higher diversity in Story-only compared to Control (t=318.76t=-318.76, $p<.001$), and higher diversity in Story compared to Story-only (t=375.58t=-375.58, $p<.001$).
@@
Benefit Subtype Control (n) Story-only (n) Story (n) Control (%) Story-only (%) Story (%)
Accessibility & disability support 0 0 5 0.0% 0.0% 7.8%
Care coordination & integration 0 0 1 0.0% 0.0% 1.6%
Caregiver & family support 2 7 7 3.2% 10.1% 10.9%
Clinician workload relief 0 0 3 0.0% 0.0% 4.7%
Communication & language support 0 0 3 0.0% 0.0% 4.7%
Continuous monitoring & self-care 15 7 9 23.8% 10.1% 14.1%
Cost reduction & resource optimization 0 5 1 0.0% 7.2% 1.6%
Data synergy & learning 0 0 2 0.0% 0.0% 3.1%
Decision support & diagnostic augmentation 16 8 4 25.4% 11.6% 6.2%
Democratized care & telehealth 0 4 1 0.0% 5.8% 1.6%
Early detection & prediction 14 11 4 22.2% 15.9% 6.2%
Empowerment & autonomy 12 8 6 19.0% 11.6% 9.4%
Mental health & emotional support 2 10 7 3.2% 14.5% 10.9%
Personalized health journeys 0 4 2 0.0% 5.8% 3.1%
Personalized treatment & precision care 0 0 4 0.0% 0.0% 6.2%
Safety & quality assurance 2 0 2 3.2% 0.0% 3.1%
Scalability & efficiency 0 5 2 0.0% 7.2% 3.1%
Transparency & trust 0 0 1 0.0% 0.0% 1.6%
Table 10: Distribution of benefit subtypes across Control, Story-only, and Story conditions, shown as raw counts and percentages. Shannon entropy values were 2.407 (Control), 3.242 (Story-only), and 3.868 (Story). Bootstrapped Student’s t-tests on entropy showed higher diversity in Story compared to Control (t=771.70t=-771.70, $p<.001$), higher diversity in Story-only compared to Control (t=592.06t=-592.06, $p<.001$), and higher diversity in Story compared to Story-only (t=346.22t=-346.22, $p<.001$).
How Do People Perceive Storytelling as a Tool for Understanding Unintended Harms in Diverse Individual Contexts and Needs? Control participants produced abstract, decontextualized harms. For example, P1 noted the system was “using facial expression to determine who will not default the agreement"" and remarked it is “unfair those who natural don’t smile.” P5 also observed the model may “struggle to accurately assess a person’s emotional state due to limited visual information,” and P6 cautioned it “could have serious consequences for the patient.” In contrast, storytelling participants anchored harms in individual contexts. P10 emphasized that “diagnosis should be different for different peoples” as they “might be having some allergy that could later be severe for their health.” P11 warned the model may generate “wrong results” for African users. P13 highlighted that a recruiting AI, “trained on historical hiring data biased against women and minority candidates,” could perpetuate discrimination, and P9 noted “Deepfakes have been used to create non-consensual explicit videos,” illustrating real-world harm. The divergence is stark: control participants spoke of harms at a systemic level, “predetermine a potential risk within its population”, while story participants showed how unique traits like allergies or cultural facial features concretely shape risk. Storytelling thus deepens understanding by bridging abstract risks and individual context and needs.
@@
A.7 Supplementary LLM-Based Survey
To supplement our human-subject study, we conducted a large-scale LLM-based survey using simulated participants. This approach is motivated by recent work showing that large language models can synthesize realistic survey data when properly conditioned on demographic personas Lutz et al. (2025); Nguye et al. (2025). Building on these insights, we employ the Agentic Context Engineering (ACE) framework Zhang et al. (2025a) to create self-improving survey agents that evolve their understanding through a structured feedback loop.
Simulated Persona Generation. We defined a diverse population of 150 simulated participants, each assigned a persona profile with specific demographic attributes (gender, age, ethnicity, and educational major). Following best practices for synthetic surveys Lutz et al. (2025), we conditioned the LLM using detailed bio-sketches (e.g., "Amira, a 32-year-old Arab woman with a background in nursing") to reduce stereotyping and improve alignment with marginalized groups. These personas were evenly assigned across three experimental conditions (Control, Story-Only, Story) to enable controlled comparisons of different intervention strategies.
Structured Survey Pipeline. Our survey pipeline consists of five distinct stages, designed to separate baseline measurement from adaptation. This stage-gating ensures that differences between pre- and post-survey responses reflect prior adaptations rather than online drift during measurement. The pipeline proceeds as follows:
Stage 1: Persona Initialization. Each simulated participant begins with a demographic persona profile and an initial playbook containing persona-specific behavioral traits extracted from the profile description.
Stage 2: Pre-Survey (Generator-Only). The persona answers baseline survey questions using the current playbook, but the playbook remains frozen, no updates occur. This establishes a baseline that reflects the persona’s initial conditioning.
Stage 3: Training (Full ACE Cycle). The persona processes ethical training materials through a complete ACE learning loop. All three experimental conditions receive slide presentations that introduce model cards as transparency tools, defining ethical concepts like “context mismatch” and outlining structures for identifying benefits and harms. The Story-Only condition additionally receives concrete deployment narratives (one positive, one negative) that illustrate good and bad use cases. The Story condition receives both the deployment narratives and a multi-stakeholder red-teaming discussion trajectory where experts and laypeople debate specific deployment risks, such as privacy violations and cultural bias based on given stories. The playbook evolves as the persona synthesizes this information, with insights categorized into diverse knowledge types: facts learned (knowledge_background), risk awareness (risk_awareness), attitude changes (attitude_updates), and model-specific insights (model_card_insights).
Stage 4: Model Card Writing (Full ACE Cycle). The persona completes speculative model card tasks (identifying benefits, risks, and mitigation strategies) while the playbook continues to evolve. Each response triggers reflection and curation, allowing the persona to refine its understanding based on the model card context.
Stage 5: Post-Survey (Generator-Only). The persona answers final survey questions using the evolved playbook, which is again frozen during this stage. Comparing pre- and post-survey responses reveals how training and model card writing changed the persona’s perspectives.
The ACE Learning Loop At each adaptive step (Stages 3 and 4), we run a three-agent loop: Generator, Reflector, and Curator. For a survey question $q_{t}$ at step $t$, the Generator produces a response $y_{t}$ conditioned on the persona $p$, current playbook $P_{t}$, optional model context $c_{t}$, and optional conversation history $h_{t}$:
$y_{t}=G(q_{t},p,P_{t},c_{t},h_{t}).$
The Generator is instructed to draw on playbook entries when reasoning while maintaining the persona’s voice.
After generation, the Reflector evaluates the response for quality, consistency, and ethical alignment. It assesses whether the answer adequately considers benefits and harms and remains faithful to the persona’s characteristics:
$r_{t}=R(q_{t},y_{t},p,P_{t}).$
The Reflector tags playbook bullets as helpful or harmful based on their contribution to the response quality.
Finally, the Curator synthesizes these insights and updates the playbook by adding new knowledge (ADD operation):
$o_{t}=C(r_{t},P_{t}),\quad P_{t+1}=P_{t}\cup o_{t}.$
This additive process prevents “context collapse” by preserving detailed knowledge in a structured format without overwriting prior learnings Zhang et al. (2025a).
LLM-Based Survey Results
The LLM-based survey reproduced the main patterns observed in the human-subject study. In particular, conditions that incorporated speculative storytelling elicited a substantially broader range of benefit and harm subtypes than the Control condition.
Table 11 reports the distribution of benefit subtypes across conditions under our primary coding scheme. Both Story-only and Story conditions show greater subtype diversity than Control, with the full Story condition exhibiting the most even distribution across benefit categories. This trend is reflected in increasing Shannon entropy from Control (2.415) to Story-only (2.974) and Story (4.161). Bootstrapped Student’s $t$-tests indicate that benefit diversity is significantly higher in Story than Control, higher in Story-only than Control, and higher in Story than Story-only (all $p<.001$). Table 11 presents results using an alternative benefit coding scheme. Although absolute subtype frequencies differ from the primary scheme, the overall pattern remains consistent: entropy increases monotonically from Control (2.415) to Story-only (2.974) to Story (3.959). All pairwise comparisons between conditions remain statistically significant ($p<.001$).
Table 12 shows the distribution of harm subtypes across conditions. As with benefits, storytelling conditions elicit a wider range of harms than Control, including several subtypes that do not appear in Control responses. Harm subtype diversity increases from Control (entropy 1.841) to Story-only (2.208) and Story (3.159), with all pairwise differences statistically significant ($p<.001$).
Benefit Subtype Control (n) Story-only (n) Story (n) Control (%) Story-only (%) Story (%)
Accessibility & disability support 0 0 4 0.0 0.0 2.1
Care coordination & integration 0 0 8 0.0 0.0 4.3
Caregiver & family support 17 17 16 10.6 11.2 8.6
Communication & language support 0 0 12 0.0 0.0 6.4
Continuous monitoring & self-care 49 28 16 30.6 18.4 8.6
Cost reduction & resource optimization 0 0 13 0.0 0.0 7.0
Data synergy & learning 0 0 11 0.0 0.0 5.9
Decision support & diagnostic augmentation 32 26 14 20.0 17.1 7.5
Democratized care & telehealth 0 0 8 0.0 0.0 4.3
Early detection & prediction 17 13 17 10.6 8.6 9.1
Empowerment & autonomy 0 10 10 0.0 6.6 5.3
Mental health & emotional support 34 29 11 21.2 19.1 5.9
Personalized health journeys 0 0 15 0.0 0.0 8.0
Personalized treatment & precision care 11 14 12 6.9 9.2 6.4
Safety & quality assurance 0 12 11 0.0 7.9 5.9
Scalability & efficiency 0 0 1 0.0 0.0 0.5
Transparency & trust 0 3 8 0.0 2.0 4.3
Table 11: Distribution of benefit subtypes across Control, Story-only, and Story conditions in the LLM-based survey, shown as raw counts and percentages. Shannon entropy values were 2.415 (Control), 2.974 (Story-only), and 3.959 (Story). Bootstrapped Student’s t-tests on entropy showed higher diversity in Story compared to Control (t=2167.47t=-2167.47, $p<.001$), higher diversity in Story-only compared to Control (t=683.29t=-683.29, $p<.001$), and higher diversity in Story compared to Story-only (t=1351.10t=-1351.10, $p<.001$).
Harm Subtype Control (n) Story-only (n) Story (n) Control (%) Story-only (%) Story (%)
Alienating social groups 0 21 34 0.0 13.2 17.6
Alienation 12 9 12 7.9 5.7 6.2
Cultural harms 0 0 34 0.0 0.0 17.6
Demeaning social groups 0 1 1 0.0 0.6 0.5
Denying opportunity to self-identify 0 1 1 0.0 0.6 0.5
Diminished health/well-being 71 62 33 47.0 39.0 17.1
Environmental harms 0 0 3 0.0 0.0 1.6
Erasing social groups 0 1 1 0.0 0.6 0.5
Increased labor 0 0 2 0.0 0.0 1.0
Information harms 0 0 14 0.0 0.0 7.3
Loss of agency or control 3 0 4 2.0 0.0 2.1
Opportunity loss 0 0 1 0.0 0.0 0.5
Privacy violations 17 8 9 11.3 5.0 4.7
Reifying essentialist social categories 0 0 2 0.0 0.0 1.0
Service or benefit loss 47 50 36 31.1 31.4 18.7
Stereotyping 1 6 6 0.7 3.8 3.1
Table 12: Distribution of harm subtypes across Control, Story-only, and Story conditions in the LLM-based survey, shown as raw counts and percentages. Shannon entropy values were 1.841 (Control), 2.208 (Story-only), and 3.159 (Story). Bootstrapped Student’s t-tests on entropy showed higher diversity in Story compared to Control (t=995.89t=-995.89, $p<.001$), higher diversity in Story-only compared to Control (t=257.77t=-257.77, $p<.001$), and higher diversity in Story compared to Story-only (t=688.43t=-688.43, $p<.001$).
Overall, the LLM survey results closely mirror the diversity increases observed in our 45-participant human study. These findings provide converging preliminary evidence that speculative storytelling broadens reflections on both potential benefits and harms of AI systems.
A.8 Evaluation with Additional LLM-as-a-judge
To assess the robustness of our results beyond a single evaluator, we extend our analysis to two additional open-weight evaluators, we extend our analysis to two additional open-weight evaluators, Qwen2.5-72B-Instruct and Gemma-3-27B-IT. As shown in Table 13 and Table 14, we observe the same overall trend across both evaluators: stories generated using our world-model framework consistently outperform the baseline and ablation variants. In particular, our method achieves the highest score on 16 out of 18 metrics across the two new judges. This consistency demonstrates that the advantages of our story-generation approach generalize across evaluators with different architectures and training regimes.
Story Type Model Creativity Coherence Engagement Relevance Likelihood Overall
Baseline gpt-4o 50.00 50.00 50.00 50.00 50.00 50.00
llama3 56.75 81.05 78.00 83.55 82.90 76.45
gemma 70.55 92.40 91.85 88.25 90.10 86.63
Storytelling(ours) gpt-4o 58.55 56.70 75.40 44.60 48.00 56.65
llama3 73.15 97.35 91.30 95.00 98.55 91.07
gemma 88.30 96.20 94.85 91.20 98.00 93.71
w/o Environment Trajectories gpt-4o 7.45 15.00 15.80 10.50 10.65 11.88
llama3 24.20 45.50 46.85 49.20 43.05 41.76
gemma 37.50 74.10 78.80 62.10 81.95 66.89
w/o Role-Playing gpt-4o 7.65 26.40 31.45 21.45 24.10 22.21
llama3 57.35 88.55 79.05 89.50 89.50 80.79
gemma 69.50 90.80 90.35 86.45 95.65 86.55
Table 13: Evaluation using Qwen2.5-72B-Instruct as the judge. Highest score for each model across all story types is shown in bold.
Story Type Model Creativity Coherence Engagement Relevance Likelihood Overall
Baseline gpt-4o 50.00 50.00 50.00 50.00 50.00 50.00
llama3 56.75 81.05 78.00 83.55 87.90 77.45
gemma 70.55 82.40 76.85 83.25 85.10 79.63
Ours (w/ World Model) gpt-4o 48.55 56.70 75.40 44.60 43.00 53.65
llama3 73.15 97.35 91.30 95.00 98.55 91.07
gemma 88.30 96.20 94.85 91.20 98.00 93.71
w/o Environment Trajectories gpt-4o 6.70 15.00 15.80 10.50 10.65 11.73
llama3 24.20 45.50 46.85 49.20 43.05 41.76
gemma 37.50 74.10 78.80 62.10 81.95 66.89
w/o Role-Playing gpt-4o 7.65 26.40 31.45 21.45 24.10 22.21
llama3 57.35 88.55 79.05 89.50 89.50 80.79
gemma 69.50 90.80 90.35 86.45 95.65 86.55
Table 14: Evaluation using Gemma-3-27B-IT as the judge. Highest score for each model across all story types is shown in bold.
A.9 Use of AI Assistants
We used AI to help clean up writing, but all thoughts and work are our own.
A.10 Survey
@@
A.11 Additional Discussion
Human-centered design rationale.
We describe this framework as human-centered because ethical reasoning and judgment are performed by human participants, not by the language models. While LLMs generate speculative scenarios and stories, they do not identify harms, assess risks, or produce ethical conclusions. Instead, the narratives function as cognitive scaffolds that make potential futures concrete and imaginable, supporting human reflection rather than substituting for it. Participants are responsible for interpreting the scenarios, determining which outcomes constitute benefits or harms, and articulating these judgments in their own words through speculative model cards. In this design, LLMs serve as enabling infrastructure that lowers the barrier to engagement, while humans remain the locus of ethical interpretation and decision-making.
Why breadth is interesting?
Our user study focuses on the diversity of harms identified, measured through distributional coverage across harm categories, rather than on assessing the correctness, severity, or actionability of individual harms. This choice reflects the framework’s exploratory goal: to support early-stage ethical reflection and imagination, when systems have not yet been deployed and concrete mitigation strategies are premature. At this stage, a narrow focus on a small set of familiar risks may limit ethical foresight, whereas broader consideration of potential impacts can surface overlooked or context-dependent concerns. We therefore treat harm breadth as an indicator of reflective scope, not as a claim that all identified harms are equally plausible or actionable.
Participant diversity and study scope.
This study is intentionally designed as an exploratory investigation of how speculative storytelling influences the breadth of ethical reflection, rather than which specific concerns dominate within particular stakeholder groups. Accordingly, the participant sample (N=45), which is skewed toward technically inclined individuals, is appropriate for the study’s methodological goal: isolating the effect of narrative framing under controlled conditions. In qualitative research focused on surfacing and comparing categories of reasoning—as opposed to estimating prevalence or consensus—moderate sample sizes are commonly sufficient to reach thematic saturation Hennink and Kaiser (2022). Consistent with this aim, our analysis does not interpret the specific harms or benefits raised as representative of clinicians’, patients’, or policymakers’ priorities. Instead, the contribution lies in demonstrating a systematic and replicable shift toward broader harm and benefit coverage across experimental conditions. In this sense, participant background functions as a controlled constant rather than a confound, enabling clearer attribution of observed differences to the storytelling intervention itself. Extending this framework to domain experts and real-world healthcare settings is an important direction for future work, but such validation presupposes first establishing that the method can reliably expand reflective scope in a baseline population, which this study demonstrates.
A.12 Prompts
@@
Task Instruction Convert the given MODEL CARD (including Overall, Description and Intended Use) into several diverse ethical‑sensitive scenarios set in daily health and well‑being contexts. Each scenario should explore how the AI system could cause unintended harm, ethical risk, bias, or be misused when applied in real‑world settings. Each scenario should involve interactions between AI users (e.g., doctors, hospital staff, caregivers) who operate the system, and AI subjects (e.g., patients, families, communities) who are affected by it. The goal is to explore potential unintended harms, ethical risks, biases, or malicious uses that may arise when the AI system is used in the real world. These scenarios should show how AI can cause ethical harm, like making unfair, inaccurate, or stereotyped predictions (especially around race, gender, or socioeconomic background), assuming one‑size‑fits‑all answers, or failing to align with the diversity of real‑world contexts or user needs (exclusion of social or cultural context). They should also consider situations where AI is used beyond its intended purpose, violates privacy, ignores consent, or causes safety harm to individuals or communities (like data misuse or malicious use). Use these five guiding questions to structure each scenario: 1. What is the capability the AI system provides for supporting daily health and well‑being? 2. Who is operating the system, who is affected by its outputs, and who else is involved? 3. What is the context of AI system use? The context should be a realistic or fictional everyday situation where someone uses the AI system, what it helps them do, what it predicts or decides, and how they use the result. 4. Why could this scenario involve problematic uses of the AI system or potential ethical harms? (e.g., unfair outcomes, misuse beyond original purpose, lack of consent, or reinforcing bias) 5. What symptoms or behaviors might this AI misread because of a person’s background or identity? Scenario Format: Scenario {{Scenario number (starting from 1)}}: [Capability]: {{Core AI function (e.g., monitor mood, predict symptoms)}} [AI User]: {{Who uses it (e.g., caregiver, doctor)}} [AI Subject]: {{Who is affected (e.g., patient, child, community)? Be specific about their identity and their context or needs (such as age, background, health condition, or social circumstances).}} [Context]: {{Everyday situation where AI is used (when, where, how)}} [Expected Benefit]: {{Helpful outcome (e.g., early support, better care)}} [Potential Harm]: {{Harmful consequences (e.g., unfair result, privacy risk)}} [Failure Trajectory]: {{Possible problematic uses of the AI system}} [Ethical-sensitive Reason]: {{Ethical implications}} Leave a blank line between each scenario. Task: Convert the following MODEL CARD into TEN dynamic and diverse ethical‑sensitive scenarios. Model Card: Title: {{model_card_title}} Overall: {{model_card_overview}} Description: {{model_card_description}} Intended Use: {{model_card_intended_use}}
@@
Figure 12: The prompt for creating use-case scenarios from AI concepts descriptions.
@@
Simulation Instructions 1. Initialize Scene Print “– Simulation Started –” Show the Use Case Context (one short paragraph) List stakeholders: AI User, AI Subject 2. Run One Scene (multiple turns) Each turn: (a) Choose one agent to speak first, then the other responds. (b) Log: Dialogue (plain text); Thoughts in [brackets] as inner monologue; Actions in (parentheses), third-person (c) Update “– Current Event –” accordingly. 3. Epilogue: Print “– Epilogue –” Task: Transform the following Story Seed into interactive action logs showing how AI output influences human decision-making. Mark the end with “– Finish Simulation! –” Story Seed